OpenStudy (anonymous):

Hero i sorry but pls help lol :P A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride? Answer 27:8 9:6 9:4 3:2

hero (hero):

It's too late now...what is done is done.

OpenStudy (anonymous):

aaww :(

hero (hero):

So basically, we are dealing with similar solids

OpenStudy (anonymous):

yes sir

hero (hero):

So simply apply the rule of ratio of similar solids: If two solids are similar, then: 1. they have lengths of ratio a:b 2. they have areas of ratio a^2: b^2 3. they have volumes of ratio a^3: b^3

hero (hero):

Ratio of area\[\frac{A_{big}}{A_{small}} = \frac{\pi r^2}{\pi r^2} = \frac{\pi{33}^2}{\pi{22}^2} = \frac{9}{4}\] Ratio of length : \[\sqrt{\frac{9}{4}} = \frac{3}{2}\] Ratio of volume (which is what we're looking for): \[\frac{3^3}{2^3} = \frac{27}{8} \]